### v3

# Stochastic Halpern iteration in normed spaces and applications to reinforcement learning

Abstract.
We analyze the oracle complexity of the stochastic Halpern iteration with variance reduction, where we aim to approximate fixed-points of nonexpansive and contractive operators in a normed finite-dimensional space. We show that if the underlying stochastic oracle is with uniformly bounded variance, our method exhibits an overall oracle complexity of $\tilde{O}(\varepsilon^{-5})$, improving recent rates established for the stochastic Krasnoselskii-Mann iteration. Also, we establish a lower bound of $\Omega(\varepsilon^{-3})$, which applies to a wide range of algorithms, including all averaged iterations even with minibatching. Using a suitable modification of our approach, we derive a $O(\varepsilon^{-2}(1-\gamma)^{-3})$ complexity bound in the case in which the operator is a $\gamma$-contraction. As an application, we propose new synchronous algorithms for average reward and discounted reward Markov decision processes. In particular, for the average reward, our method improves on the best-known sample complexity.
Key words and phrases:
stochastic fixed-point iterations, nonexpansive maps, error bounds, convergence rates, $Q$-learning
2010 Mathematics Subject Classification:
47J25, 47J26, 65J15, 65K15
Acknowledgements. Mario Bravo gratefully acknowledges the support by FONDECYT grant 1191924 and the Anillo grant ACT210005. Juan Pablo Contreras was supported by Postdoctoral Fondecyt Grant No. 3240505.
1. Introduction
Fixed-point iterations are fundamental in mathematics and engineering, providing a systematic way to approximate the solution of several problems, including operator equations, nonlinear equations, and optimization problems. These iterations form the backbone of many optimization algorithms, such as the gradient descent or splitting method, which are widely used in machine learning, control systems, and scientific computing.
In recent years, stochastic variants of these problems have gained significant attention because of their applicability in scenarios where the underlying operators are not known exactly or are subject to noise, uncertainty, or randomness. Nevertheless, general stochastic fixed-point iterations have received little attention compared to other related setups, such as stochastic monotone inclusions, stochastic variational inequalities, or stochastic optimization which are mainly developed in the Euclidean setting.
The main goal of this work is then to study a simple stochastic version of the well-known Halpern iteration, where we assume that the underlying operator is nonexpansive or contracting for some general norm in finite dimension. Using a simple approach derived mainly from the construction of recursive bounds and a simple variance reduction technique, we study the oracle complexity of this algorithm while showing applications that are intrinsically of a non-Euclidean nature.
1.1. Problem setup
[…]
$\left\|Tx-Ty\right\|\leq\gamma\left\|x-y\right\|,\qquad\forall x,y\in{\mathbb{R}}^{d}.$
We say that $T$ is a contraction if $\gamma\in(0,1)$, and that $T$ is nonexpansive if $\gamma=1$. We assume that in the nonexpansive case $\mathop{\rm Fix}\,T\neq\emptyset$, where $\mathop{\rm Fix}\,T$ stands for the set of fixed-points of $T$. We are interested in computing an approximate solution of the fixed-point equation $x=Tx$.
Our method is inspired by the standard Halpern iteration
(H) $x^{n}=(1-\beta_{n})x^{0}+\beta_{n}Tx^{n-1},\quad\forall\,n\geq 1,$
where $(\beta_{n})_{n}$ is a nondecreasing stepsize sequence satisfying that $10<\beta_{n}<$ 1, $\lim_{n\to\infty}\beta_{n}=1$, and $x^{0}\in{\mathbb{R}}^{d}$ is an arbitrary initial condition. It is worth mentioning that the original Halpern iteration uses an arbitrary $u$ called the anchoring point that can be different from the initial point $x^{0}$. In that case, the iteration reads $x^{n}=(1-\beta_{n})u+\beta_{n}Tx^{n-1}$. We assume here, without loss of generality, that the anchoring point is exactly $x^{0}$.
Naturally, iteration (H) is only implementable in the case where one has access to $Tx$ at any point $x$. In this paper, we assume that the vector $Tx$ is only accessible through $\tilde{T}$ which is a noisy version of $T$. Specifically, we consider the following
Assumption 1.1.
The access to the operator $T$ occurs through a stochastic oracle $\tilde{T}:{\mathbb{R}}^{d}\times\Xi\to{\mathbb{R}}^{d}$ which is an unbiased approximation of $T$ in the sense that
${\mathbb{E}}(\tilde{T}(x,\xi))=Tx\qquad\mbox{for all $x\in{\mathbb{R}}^{d}$},$
with $\xi$ drawn from some distribution $\mathcal{D}_{x}$ over $\Xi$. Additionally, we assume that ${\mathbb{E}}\left(\left\|\tilde{T}(x,\xi)-Tx\right\|^{2}_{2}\right)$ is finite for every $x\in{\mathbb{R}}^{d}$.
We say that the oracle $\tilde{T}$ is with uniformly bounded variance (with bound $\sigma^{2})$ if there exists a positive constant $\sigma^{2}$ such that
$\sup_{x\in{\mathbb{R}}^{d}}{\mathbb{E}}\left(\left\|\tilde{T}(x,\xi)-Tx\right\|^{2}_{2}\right)\leq\sigma^{2}.$
A first possible idea is to consider the iteration $x^{n}=(1-\beta_{n})x^{0}+\beta_{n}(\tilde{T}x^{n-1})$. Intuitively, even if the term $\tilde{T}(x,\xi)-Tx$ is with uniformly bounded variance, because $\beta_{n}$ goes to 1, the last iterate will inevitably contain some nonnegligible noise term. To address this issue, we adopt a simple minibatch strategy designed to mitigate the variance of the error. This approach lead to the study of the following inexact version of the Halpern iteration (H),
(iH) $x^{n}=(1-\beta_{n})x^{0}+\beta_{n}(Tx^{n-1}+U_{n}),\quad\forall\,n\geq 1,$
where $U_{n}$ is such that ${\mathbb{E}}(U_{n})=0$ and the term ${\mathbb{E}}(\left\|U_{n}\right\|)$ can be controlled appropriately. It is worth noting that we do not assume further properties on the stochastic oracle in order to maintain our results as general as possible. See Section 2 for a precise definition and related discussions.
We are interested in the problem of estimating the oracle complexity to find and approximate fixed-point of $T$, which is measured as the number of calls to the vector stochastic oracle $\tilde{T}$. In the case where $T$ is nonexpansive, we estimate the oracle complexity to find a last-iterate approximate solution $x^{n}$ in the sense that the expected fixed-point residual ${\mathbb{E}}(\left\|x^{n}-Tx^{n}\right\|)$ is smaller than a given tolerance. For a contracting operator $T$, we aim to find oracle complexity bounds for the expected distance of the last iterate $x^{n}$ to the unique fixed-point $x^{*}$, i.e. the quantity ${\mathbb{E}}(\left\|x^{n}-x^{*}\right\|)$.
$\bullet$ Notation. Let us set some basic notation that we use in this paper. Given that all norms in ${\mathbb{R}}^{d}$ are equivalent, let $\mu>0$ be such that $\left\|x\right\|\leq\mu\left\|x\right\|_{2}$ for all $x\in{\mathbb{R}}^{d}$. We adopt the standard big-$O$ notation, that is, for two real functions $f,g:$ we say $f(x)=O(g(x))$ if there are two constants $C$ and $c$ such that $f(x)\leq Cg(x)$ for all $x\geq c$. We use $\tilde{O}$ if we ignore logarithmic factors. Analogously, $f(x)=\Omega(g(x))$ if $f(x)\geq Cg(x)$ for all $x\geq c$.
1.2. Our contributions
In what follows, we briefly describe our main findings.
$\bullet$ Nonexpansive maps. We show that if $T$ is nonexpansive and the stochastic oracle is with bounded variance, then our method achieves last-iterate oracle complexity of $\tilde{O}(\varepsilon^{-5})$ (see Corollary 3.5). This result improves over the guarantee $\tilde{O}(\varepsilon^{-6})$ established for the stochastic Krasnoselskii-Mann iteration in [15] also for the last-iterate. This is, to the best of our knowledge, the only complexity result available at the moment in this general framework.
Additionally, we establish a lower bound of $\Omega(\varepsilon^{-3})$ for the overall oracle complexity for nonexpansive maps with uniformly bounded variance stochastic oracle and where the underlying operator has bounded range. Our result is also true in the case where the use of minibatches is allowed (c.f. Theorem 3.7). In particular, this lower bound applies to every algorithm such that the iterations are a linear combination of the previous (possibly variance-reduced) oracle calls, including well-known algorithms such as the Krasnoselskii-Mann and Halpern iteration.
Finally, we extend the results in [54] by proving that the iteration converges to a fixed point of $T$ provided that the norm of the space is smooth (c.f. Theorem 3.10).
$\bullet$ Contractions. In the case where $T$ is a contraction with parameter $\gamma\in(0,1)$, we prove that if the stochastic oracle is with uniformly bounded variance, then our method finds an $\varepsilon-$optimal solution after at most $O\left(\varepsilon^{-2}(1-\gamma)^{-3}\right)$ queries (Corollary 4.3). Given that this is a general result, the bound depends on the specific instance and can be refined if more structure is available. Indeed, the same holds for the nonexpansive case.
Nevertheless, it is worth noting that, as pointed out in recent work [34], little attention has been devoted to the study of stochastic fixed-point iterations for contracting operators without assuming some form of monotonicity for $T$.
$\bullet$ Markovian Decision Processes. First, we apply our analysis to the problem of a Markovian decision process (MDP) with average reward. Arguably, this is the most relevant setting for reinforcement learning, where the results are much less complete compared to the case of discounted rewards [51, 45]. In Theorem 5.1, we show that a suitable modification of our main method results in a new synchronous algorithm with sample complexity $\tilde{O}\left(\varepsilon^{-7}\right)$. To our knowledge, this is currently the best convergence guarantee available for this problem, improving the ${O}\left(\varepsilon^{-a}\right)$, $a>10$, complexity (see [15]) recently obtained while analyzing the RVI-$Q$- learning algorithm [1]. The most interesting aspect of our method is that it constitutes one prominent example where the underlying stochastic oracle is not necessarily with uniformly bounded variance, since no a priori uniform bound is known on the state variable.
Next, we extend our results to synchronous discounted-reward $Q$-learning, a well-researched problem. Our new algorithm offers competitive convergence guarantees and reaches a tolerance of $\varepsilon>0$ after at most $\tilde{O}\left(\varepsilon^{-2}(1-\gamma)^{-5}\right)$ samples (see Theorem 5.2), which matches recent, yet suboptimal, results in the literature [50]. Despite its suboptimal rate, our simple analysis demonstrates the effectiveness of alternative iterations in addressing the problem and encourages further research to discover superior methods. Furthermore, as in the average reward framework, our argument makes no use of the monotonicity properties of the underlying Bellman operator.
$\bullet$ Our Approach. The techniques we use to obtain our results are mainly based on a recursive bound approach developed first to the study of the well-known Krasnoselskii-Mann iteration,
(KM) $x^{n}=(1\!-\!\alpha_{n})x^{n}+\alpha_{n}Tx^{n-1}\quad n\geq 1$
in Banach spaces, where $(\alpha_{n})_{n}$ is a sequence of stepsizes in $(0,1)$ [8, 18, 14]. Later, the analysis was extended to general averaged fixed-point iteration [13]. In the case of the Halpern iteration, being a particular case of an averaged iteration, the bounds developed are particularly explicit, giving way to an analysis that is simple yet effective. Roughly speaking, in the stochastic setting, we will adapt these ideas by compensating the effect of repeated calls to the oracle with the smooth contracting effect of the term $\beta_{n}<1$. Given that $\beta_{n}$ approaches one, we will take increasing minibatch sizes. It is worth mentioning that our explicit results are obtained for the case $\beta_{n}=\frac{n}{n+1}$. The reason behind this particular choice is twofold. On the one hand, it is known that in the noiseless case this stepsize gives the best possible bounds in the Euclidean setting [44, 33] and, up to log factors, also in the Banach framework. On the other hand, as we shall see in what follows, this choice leads to clean estimates with computable constants when calculating fixed-point residuals.
1.3. Related work
$\bullet$ Deterministic setup. For nonexpansive maps in general normed spaces, recent efforts have been devoted to establish nonasymptotic convergence guarantees for the fixed-point residual $\left\|x^{n}-Tx^{n}\right\|$ of some classical iterations. For (KM), convergence rates were initially determined by [7, 8] for constant stepsizes $\alpha_{n}\equiv\alpha$, showing a rate of $O(1/\sqrt{n})$. Later, these results were extended by [18] to arbitrary steps in $(0,1)$. The sharpness of the bound in [18] was obtained in [14]. The first proof of the rate $O(1/n)$ was established in [44] using the Halpern iteration (H) with the specific stepsize $\beta_{n}=\frac{n}{n+2}$. Later, in [13], it was shown that the traditional Halpern stepsize choice $\beta_{n}=\frac{n}{n+1}$ only yields the (tight) rate $O(\frac{\log n}{n})$, while in [19] the optimal coefficients for the Halpern iteration were determined, improving the constants involved.
For Euclidean spaces, the exact upper bounds were determined using the Performance Estimation Problems (PEP) approach, originally proposed in [23]. However, the convergence rate does not show an improvement compared to the general Banach setting. The classical Halpern iteration with stepsize $\beta_{n}=\frac{n}{n+1}$ was examined in [33] and the authors proved the convergence rate $O(1/n)$, which improves by a logarithmic factor compared to the convergence rate in general normed spaces.
Halpern iteration has also been successfully applied in the context of monotone inclusions, which is strongly connected to fixed-point iterations [9]. In particular, using PEP, [29] determined an optimal inertial iteration to approximate the zeros of a cocoercive operator that was shown to be equivalent to the Halpern iteration in [19] and [43, Chapter 12.2, Theorem 18]. Beyond PEP, convergence rates have also been established using a potential function approach [20, 21, 47, 30]. It is also known that, in convex optimization, Halpern iteration is connected to well-established first-order methods such as Nesterov’s accelerated gradient descent [46, 36] and extra-gradient type methods [48, 39].
$\bullet$ Stochastic setup. In the Euclidean setting, Halpern iteration was studied to approximate the zeros of a cocoercive operator in [16]. Using variance reduction based on minibatches of quadratic growth, combined with ideas developed for the deterministic setting [20], the authors established an overall oracle complexity of $\tilde{O}(\varepsilon^{-4})$ to find an approximate solution in expectation within a tolerance $\varepsilon$. The authors also show that the rate $\tilde{O}(\varepsilon^{-3})$ is achievable under further assumptions on the stochastic oracle, such as the multi-point oracle and Lipschitz in expectation properties (c.f. Section 2 for details). Under the same enhanced stochastic oracle, [32] obtained similar results in the cocoercive setting. Very recently, [3] studied a stochastic Halpern iteration for cohypomonotone operators [10], a condition that allows possible nonmonotonicity while including monotone maps. The authors establish, under no further assumptions on the oracle structure, a complexity of $\tilde{O}(\varepsilon^{-4})$ for the fixed-point residual.
Up to our knowledge, the only work that studies convergence rates for stochastic fixed-point iterations of nonexpansive maps in general normed spaces is [15]. Here, the authors consider the stochastic Krasnoselskii-Mann iteration $x^{n}=(1\!-\!\alpha_{n})x^{n}+\alpha_{n}(Tx^{n-1}+U_{n})$, where $U_{n}$ is assumed to be a martingale difference sequence. The authors prove that if the variance of the noise is uniformly bounded, the stochastic Krasnoselskii-Mann iteration with constant stepsize requires at most $O(\varepsilon^{-6})$ iterations to reach a solution whose expected fixed-point residual is less than $\varepsilon$. This is done using a well-chosen constant stepsize.
$\bullet$ Variance reduction. Variance reduction methods are effective techniques to improve convergence rates in stochastic settings and have been mainly developed for stochastic optimization. In [28] the authors propose the SVRG method for convex optimization problems with a finite-sum structure. Later, [4] extended the technique to find stationary points of nonconvex functions. Using recursive and nested variance reduction methods [25] and [55] improved the convergence rate in the context of nonconvex sums. Later, in [31] the authors obtained a simplified algorithm that preserves the optimal rate and extended the result beyond the small data regime of [25] and [55]. For this, they used variance reduction based on a probabilistic gradient estimator (PAGE). In fact, a suitable modification of PAGE is an important tool for the results obtained for the stochastic Halpern iteration [16].
Beyond optimization, variance reduction methods have been successfully implemented in the setting of stochastic variational inequalities. In their work [37], the authors expanded the SVRG algorithm to address convex-concave saddle-point problems incorporating splitable functions. In [27] proposed an extragradient method with a variance reduction method based on minibatch. Their algorithm has near-optimal oracle complexity, and covers the important case of unbounded operators and unbounded feasible sets. Similarly, in [17] a new algorithm that combines variance reduction and extrapolation is studied in order to achieve more stable training in setting of generative adversarial networks.
$\bullet$ Lower bounds. In the noiseless scenario, the lower bound for fixed-point iterations was determined in [19] showing that the rate $\left\|x^{n}-Tx^{n}\right\|=O(1/n)$ is optimal for every averaged iteration. Moreover, this rate cannot be improved even in spaces with more structure, such as Hilbert spaces. In particular, [38] proved that no method such that $x^{n}\in\mbox{span}\{x^{0},Tx^{0},\dots,Tx^{n-1}\}$ can achieve a better convergence rate.
For stochastic convex optimization, the results date back to the information-based complexity developed in [35]. In this line, improvements in the dimension dependence were obtained in [2] for the setup of statistical estimation. Similar results were obtained in [41] with a different technique that yields lower bounds for a special class of descend methods that are not covered in [2]. Finding stationary points of convex functions relates to finding fixed-points of nonexpansive operators. In this context, the work [26] established the lower bounds for stochastic convex optimization problems in the local and global stochastic oracle model. For nonconvex problems, the lower bound was established for the first time in [25] for the sum of smooth functions, matching the upper bounds of the variance-reduced methods SPIDER [25] and SNVRG [55]. For further results and discussion, we refer the reader to [22, 5] and the references therein.
$\bullet$ Synchronous $Q$-learning. Given its wide range of application, discounted reward $Q$-learning has been extensively studied in the literature [40, 45, 53, 52]. In fact, the synchronous version of the vanilla $Q$-learning algorithm corresponds exactly to a stochastic version of the Krasnoselskii–Mann iteration, where the operator involved is contracting.
Oracle complexity bounds have been obtained for constant [11], linear polynomial [24], and linearly rescaled learning rates [49, 50]. For the model-based version where, roughly speaking, transitions are sampled previously with a full batch and then a value iteration is performed, in [6] the authors obtained a complexity bound of $O(\epsilon^{-2}(1-\gamma)^{-3})$. Moreover, they show that it is indeed minimax optimal. For the classical algorithm, [49] obtained the rate $\tilde{O}(\epsilon^{-2}(1-\gamma)^{-5})$. Later, [50] reached the optimal complexity bound, up to logarithmic factors, using a variance reduction strategy. See [50] for a full history on this matter.
On the other hand, $Q$-learning with average reward has received much less attention [45, 51]. In fact, the sample complexity bound for one of the few algorithms available, the RVI $Q$-learning algorithm [1] was established only recently in [15].
1.4. Organization of the paper
The remainder of the paper is organized as follows. In Section 2 we present the precise definition of our method along with the minibatch variance reduction. Section 3 is devoted to the study of the case in which $T$ is nonexpansive while in Section 4 we study the contracting case. In Section 5 we develop our applications to MDPs in the average and discounted setting. Finally, Appendix A gathers some deferred proofs.
2. The Stochastic Halpern iteration
Algorithm 1 outlines the stochastic Halpern iteration with minibatches. At each iteration $n\leq N$, a minibatch estimator is constructed with $k_{n}$ queries to $\tilde{T}$ at $x^{n-1}$ using $k_{n}$ independent fresh samples from $\mathcal{D}_{x^{n-1}}$. Using this estimate, the next iterate is calculated via the Halpern iteration.
Algorithm 1 Stochastic Halpern method
1:Stochastic oracle $\tilde{T}$, $(k_{n})_{n}$ minibatch size sequence, number of iterations $N$, initial point $x^{0}$
2:for $n=1,\dots,N$ do
3:     Set the size of the minibatch $k_{n}$
4:     Sample $k_{n}$ independent queries at $x^{n-1}$: $\tilde{T}(x^{n-1},\xi_{n,1}),\dots,\tilde{T}(x^{n-1},\xi_{n,k_{n}})$.
5:     Set $x^{n}\leftarrow(1-\beta_{n})x^{0}+\beta_{n}\left(\frac{1}{k_{n}}\displaystyle{\sum_{j=1}^{k_{n}}}\tilde{T}(x^{n-1},\xi_{n,j})\right)$
6:end for
7:Return $x^{N}$
Notice that if we define the error term at the iteration $n\geq 1$ by
$U_{n}:=\frac{1}{k_{n}}\displaystyle{\sum_{j=1}^{k_{n}}}\tilde{T}(x^{n-1},\xi_{n,j})-Tx^{n-1},$
then Algorithm 1 can be cast in the form (iH). Indeed, we will refer to both procedures indistinctly when there is no confusion.
By construction, $Tx^{n-1}+U_{n}$ is an unbiased estimator of $Tx^{n-1}$ and may have a reduced variance compared to the error computed with a single query. Let us introduce the following additional notation: for all $i,n\in{\mathbb{N}}$
(1) $\sigma_{n}:={\mathbb{E}}(\left\|U_{n}\right\|),\quad\text{and}\quad B_{i}^{n}:=\prod_{j=i}^{n}\beta_{j},$
with the convention that $\sigma_{0}=\beta_{0}=0$ and $B_{i}^{n}=1$ if $i>n\geq 0$.
For the sake of completeness, we recall how a standard argument allows one to transfer the variance reduction into a bound on $\sigma_{n}$. For a given iteration $n$, let $\xi_{n,1},\ldots,\xi_{n,k_{n}}$ be $k_{n}$ independent samples of the distribution $\mathcal{D}_{x^{n-1}}$ and let $e_{n}^{2}$ be the variance of the error term of one query $\tilde{T}(x^{n-1},\xi_{n,1})-Tx^{n-1}$. Then
(2) $\sigma_{n}^{2}={\mathbb{E}}(\left\|U_{n}\right\|)^{2}\leq\frac{\mu^{2}}{k_{n}^{2}}{\mathbb{E}}\left(\left\|\sum_{j=1}^{k_{n}}\tilde{T}(x^{n-1},\xi_{n,j})-Tx^{n-1}\right\|^{2}_{2}\right)\leq\frac{\mu^{2}}{k_{n}}e_{n}^{2},$
where we have use Jensen’s inequality, sample independence, and the equivalence between the norms. We observe that ${\mathbb{E}}(\left\|U_{n}\right\|)$ scales by a factor $\frac{\mu}{\sqrt{k_{n}}}$ compared to $e_{n}$.
As we already mentioned, there exist more sophisticated variance reduction techniques to control $\sigma_{n}$. However, they often necessitate additional assumptions regarding the stochastic oracle’s nature or assume extra operator structure.
For example, in the Euclidean framework, recursive variance reduction has been employed to enhance the oracle complexity of the stochastic Halpern iteration. In [16], this entails reliance on a multi-point oracle, capable of assessing different points $x^{1},x^{2},\dots,x^{m}$ using the same random seed $\xi$. Additionally, they assume a Lipschitz in expectation property, linked to the mean-square smoothness in nonconvex optimization [25, 55]. In the realm of nonexpansive fixed-point iterations, this translates to nonexpansiveness in expectation, that is,
(3) ${\mathbb{E}}\left(\left\|\tilde{T}(x,\xi)-\tilde{T}(y,\xi)\right\|^{2}_{2}\right)\leq\left\|x-y\right\|^{2}_{2},\quad\forall x,y\in{\mathbb{R}}^{d}.$
Despite the advances made possible by these advanced variance reduction techniques, their additional hypotheses limit the scope of the analysis. Hence, we opt for the simplest variance reduction method: minibatches. By doing so, we intentionally refrain from imposing extra assumptions on the stochastic oracle, thereby ensuring the broad applicability of our analysis.
3. The nonexpansive case
In this section, we assume that $T$ is a nonexpansive operator. As we mentioned, we are interested in the number of iterations and the overall oracle complexity required by iteration (iH) to obtain a solution $x^{N}$ such that ${\mathbb{E}}(\left\|x^{N}-Tx^{N}\right\|)\leq\varepsilon$ for a given tolerance $\varepsilon>0$. The following condition will be crucial for the results of this section, since it allows to write some recursive bound for the expected fixed-point residual.
($H_{1}$) There exists $\text{There exists }\overline{\kappa}<+\infty\,\text{ such that }\,\sup\nolimits_{n\geq 0}{\mathbb{E}}(\left\|Tx^{n}-x^{0}\right\|)\leq\overline{\kappa}<+\infty.$
3.1. Bounds for the expected fixed-point residual
The following proposition provides a bound for ${\mathbb{E}}(\left\|x^{n}-Tx^{n}\right\|)$, valid for any increasing sequence of steps $(\beta_{n})_{n}\subseteq(0,1)$.
Proposition 3.1.
Let $(x^{n})_{n\geq 1}$ the sequence generated by the Halpern iteration (iH). Assume that ($H_{1}$) holds for some $\overline{\kappa}\geq 0$. Then, for all $n\geq 1$
${\mathbb{E}}(\left\|x^{n}-Tx^{n}\right\|)\leq\overline{\kappa}(1-\beta_{n})+\sum_{i=1}^{n}B_{i}^{n}\left(\overline{\kappa}(\beta_{i}-\beta_{i-1})+\beta_{i}\sigma_{i}+\beta_{i-1}\sigma_{i-1}\right)+\beta_{n}\sigma_{n}.$
Proof.
Recall that by convention $\sigma_{0}=\beta_{0}=0$ and let $\kappa_{n}=\left\|Tx^{n}-x^{0}\right\|$. We start by observing that the (random) sequence $(P_{n})_{n}$ defined inductively as $P_{0}=0$ and
$P_{n}=\kappa_{n-1}(\beta_{n}-\beta_{n-1})+\beta_{n-1}P_{n-1}+\beta_{n}\left\|U_{n}\right\|+\beta_{n-1}\left\|U_{n-1}\right\|,\,\,n\geq 1,$
is such that $\left\|x^{n}-x^{n-1}\right\|\leq P_{n}$ for all $n\in{\mathbb{N}}$. Indeed, for the first term we write
$\left\|x^{1}-x^{0}\right\|=\beta_{1}\left\|x^{0}-Tx^{0}+U_{1}\right\|\leq\kappa_{0}(\beta_{1}-\beta_{0})+\beta_{1}\left\|U_{1}\right\|=P_{1}.$
From the definition of (iH), using again the triangular inequality, the nonexpansivity of $T$, and the induction hypothesis, we have
$\!\!\left\|x^{n}\!-\!x^{n-1}\right\|$ $=\left\|(1\!-\!\beta_{n})x^{0}+\beta_{n}(Tx^{n-1}+U_{n})-(1\!-\!\beta_{n-1})x^{0}\!-\!\beta_{n-1}(Tx^{n\!-\!2}+U_{n\!-\!1})\right\|$
$=\left\|(\beta_{n}\!-\!\beta_{n-1})(Tx^{n-1}\!-\!x^{0})\!+\!\beta_{n-1}(Tx^{n-1}\!\!-Tx^{n-2})\!+\!\beta_{n}U_{n}\!-\!\beta_{n-1}U_{n-1}\right\|$
$\leq(\beta_{n}-\beta_{n-1})\left\|Tx^{n-1}\!-\!x^{0}\right\|\!+\!\beta_{n-1}\left\|x^{n-1}\!-\!x^{n-2}\right\|\!+\!\beta_{n}\left\|U_{n}\right\|\!+\!\beta_{n-1}\left\|U_{n-1}\right\|$
$\leq\kappa_{n-1}(\beta_{n}-\beta_{n-1})+\beta_{n-1}P_{n-1}+\beta_{n}\left\|U_{n}\right\|\!+\!\beta_{n-1}\left\|U_{n-1}\right\|=P_{n}.$
Then, it follows that $P_{n}$ can be computed explicitly as
$P_{n}=\sum_{i=1}^{n}B_{i}^{n-1}\left(\kappa_{i-1}(\beta_{i}-\beta_{i-1})+\beta_{i}\left\|U_{i}\right\|+\beta_{i-1}\left\|U_{i-1}\right\|\right).$
Now, casting (iH) as
$x^{n}-Tx^{n}=(1-\beta_{n})(x^{0}-Tx^{n})+\beta_{n}(Tx^{n-1}-Tx^{n}+U_{n}),\quad\forall\,n\geq 1,$
it follows that
(4) $\!\!\!\left\|x^{n}\!\!-\!\!Tx^{n}\right\|$ $\!\!=\!\!\left\|(1\!-\!\beta_{n})(x^{0}-Tx^{n})+\beta_{n}(Tx^{n-1}-Tx^{n}+U_{n})\right\|$
$\leq\kappa_{n}(1-\beta_{n})+\beta_{n}\left\|x^{n}-x^{n-1}\right\|+\beta_{n}\left\|U_{n}\right\|$
$\leq\kappa_{n}(1-\beta_{n})+\beta_{n}P_{n}+\beta_{n}\left\|U_{n}\right\|,$
$=\kappa_{n}(1-\!\beta_{n})\!+\!\sum_{i=1}^{n}B_{i}^{n}\left(\!\kappa_{i-1}(\beta_{i}-\beta_{i-1})\!+\!\beta_{i}\left\|U_{i}\right\|\!\!+\!\!\beta_{i\!-\!1}\left\|U_{i\!-\!1}\right\|\!\right)\!+\!\beta_{n}\left\|U_{n}\right\|.$
Taking expectations in the previous expression, we obtain
${\mathbb{E}}(\left\|x^{n}\!\!-\!\!Tx^{n}\right\|)$ $\leq{\mathbb{E}}(\kappa_{n})(1-\!\beta_{n})+\sum_{i=1}^{n}B_{i}^{n}\left({\mathbb{E}}(\kappa_{i-1})(\beta_{i}-\beta_{i-1})\!+\!\beta_{i}\sigma_{i}+\beta_{i\!-\!1}\sigma_{i-1}\!\right)\!+\!\beta_{n}\sigma_{n}.$
The conclusion follows from the fact that $\sup_{n\geq 0}{\mathbb{E}}(\kappa_{n})\leq\overline{\kappa}$. ∎
We complement the result in Proposition 3.1 with the following lemma, which provides two simple independent and sufficient conditions for ($H_{1}$) to hold.
Lemma 3.2.
Let $T$ be nonexpansive and $(x^{n})_{n}$ be the sequence generated by (iH). Recalling the notation (1), it holds:
  • $(i)$
    If $T$ has bounded range i.e. there exist a constant $M>0$ such that $\left\|Tx\right\|\leq M$ for all $x\in{\mathbb{R}}^{d}$, then the condition ($H_{1}$) holds with $\bar{\kappa}=M+\left\|x^{0}\right\|$.
  • $(ii)$
    If the sequences $(\beta_{n})_{n}$ and $(\sigma_{n})_{n}$ satisfies
    ($H_{2}$) $S:=\sup_{n\geq 1}\sum_{i=1}^{n}B_{i}^{n}\sigma_{i}<+\infty,$
    then condition ($H_{1}$) holds with $\bar{\kappa}=2\operatorname{dist}(x^{0},\mathop{\rm Fix}(T))+{S}$.
    In particular, if $\sum_{n}\sigma_{n}<+\infty$ then ($H_{1}$) holds.
Proof.
Part $(i)$ follows from the triangular inequality. For part $(ii)$, let us fix $x^{*}\in\mathop{\rm Fix}(T)$. Using the fact that $T$ is nonexpansive we have
$\left\|x^{n}-x^{*}\right\|$ $=\left\|(1-\beta_{n})(x^{0}-x^{*})+\beta_{n}(Tx^{n-1}-x^{*})+\beta_{n}U_{n}\right\|$
$\leq(1-\beta_{n})\left\|x^{0}-x^{*}\right\|+\beta_{n}\left\|x^{n-1}-x^{*}\right\|+\beta_{n}\left\|U_{n}\right\|$
$=\sum_{i=0}^{n}B^{n}_{i+1}(\left\|x^{0}-x^{*}\right\|(1-\beta_{i})+\beta_{i}\left\|U_{n}\right\|)$
(5) $=\left\|x^{0}-x^{*}\right\|+\sum_{i=1}^{n}B^{n}_{i}\left\|U_{i}\right\|,$
where we used $\sum_{i=0}^{n}B^{n}_{i+1}(1-\beta_{i})=1$ and the convention $B^{n}_{n+1}=1$. Taking expectations, we have
${\mathbb{E}}(\left\|Tx^{n}-x^{0}\right\|)$ $\leq{\mathbb{E}}(\left\|Tx^{n}-Tx^{*}\right\|)+\left\|x^{*}-x^{0}\right\|$
$\leq 2\left\|x^{*}-x^{0}\right\|+\sum_{i=1}^{n}B_{i}^{n}\sigma_{i}\leq 2\left\|x^{*}-x^{0}\right\|+S.$
Finally,
$\sup_{n\geq 0}{\mathbb{E}}(\left\|Tx^{n}-x^{0}\right\|)\leq 2\inf_{x^{*}\in\mathop{\rm Fix}T}\left\|x^{*}-x^{0}\right\|+S.$
The following result establishes the convergence rate for the expected residual of the Halpern iteration in the special case where $\beta_{n}=\frac{n}{n+1}$. This stepsize is easy to handle and, more importantly, leads to optimal residual estimations in the noiseless case in the Hilbert setting, and near-optimal for general Banach spaces. Moreover, it is indeed possible to write a recursive scheme for the optimal choice of the stepsize $\beta_{n}$ as in the noiseless case (see [19, Theorem 2] for precise details). Numerical experiences seem to show that the bounds improve only by a logarithmic factor. We left this intriguing question for future research.
Theorem 3.3.
Let $(x^{n})_{n}$ be the sequence generated by the Halpern iteration (iH) with the stepsize $\beta_{n}=\frac{n}{n+1}$. Assume that $\sigma_{n}={\mathbb{E}}(\left\|U_{n}\right\|)$ is such that $\sum_{n=1}^{N}n\sigma_{n}=\tilde{O}(1)$. Then, given $\varepsilon>0$, ${\mathbb{E}}(\left\|x^{N}-Tx^{N}\right\|)\leq\varepsilon$ after $N=\tilde{O}(\varepsilon^{-1})$ iterations.
Proof.
Because $\sum_{n=1}^{N}n\sigma_{n}=\tilde{O}(1)$ we have $S=\sup_{N\geq 1}\frac{1}{N+1}\sum_{n=1}^{N}n\sigma_{n}<\infty$, and therefore condition ($H_{2}$) holds. Now, invoking the estimates given by Proposition 3.1 and taking expectation we have
(6) ${\mathbb{E}}(\left\|x^{N}-Tx^{N}\right\|)$ $\leq\frac{\bar{\kappa}}{N+1}+\sum_{n=1}^{N}\frac{n}{N+1}\left(\frac{\bar{\kappa}}{n(n+1)}+\frac{2n+1}{n+1}\sigma_{n}\right)$
$\leq\frac{1}{N+1}\left(\bar{\kappa}+\bar{\kappa}\sum_{n=1}^{N}\frac{1}{n+1}+2\sum_{n=1}^{N}n\sigma_{n}\right)$
We observe that the term in the brackets is $\tilde{O}(1)$. Therefore, ${\mathbb{E}}(\left\|x^{N}-Tx^{N}\right\|)=\tilde{O}(1/N)$, and ${\mathbb{E}}(\left\|x^{N}-Tx^{N}\right\|)\leq\varepsilon$, after at most $N=\tilde{O}(\varepsilon^{-1})$ iterations as claimed. ∎
Remark 3.4.
Two comments are in order.
$\bullet$ Similar convergence rates have been proved for the deterministic Halpern iteration in Hilbert [33], and in Banach spaces [13]. Also, for the stochastic iteration in [16], and for the inexact iteration in [32]. To the best of our knowledge, our result is the first that covers the case of general normed spaces in stochastic and inexact setups.
$\bullet$ The convergence rate above is expressed in terms of the number of iterations. It is important to note that the number of queries to the stochastic oracle may not be proportional to the number of iterations due to the use of variance reduction.
The following corollary gives an upper bound on the number of queries for the stochastic Halpern iteration with minibatch when the variance is uniformly bounded. Notice that in this case the sequence $(\sigma_{n})_{n}$ is summable.
Corollary 3.5.
Assume that $\tilde{T}$ has uniformly bounded variance with constant $\sigma^{2}$. Then, given $\varepsilon>0$, Algorithm 1 with $k_{n}=n^{4}$ returns a solution $x^{N}$ such that ${\mathbb{E}}(\left\|x^{N}-Tx^{N}\right\|)\leq\varepsilon$ using at most $\tilde{O}(\varepsilon^{-5})$ queries to the stochastic oracle.
Proof.
Equation (2) guarantees that $\sigma_{n}\leq\frac{\mu\sigma}{n^{2}},$ which implies that the hypotheses of Theorem 3.3 are satisfied. Hence, ${\mathbb{E}}(\left\|x^{N}-Tx^{N}\right\|)\leq\varepsilon$ after $N=\tilde{O}(\varepsilon^{-1})$ iterations. As each iteration $n=1,...,N$ requires $k_{n}$ queries to the stochastic oracle, we find an overall oracle complexity of $\sum_{n=1}^{N}k_{n}=\sum_{n=1}^{N}n^{4}\leq N^{5}/5$. Now, retaking the bound (6) given in the proof of Theorem 3.3 and using some simple estimates, we have that the $N$-th iteration satisfies
${\mathbb{E}}(\left\|x^{N}-Tx^{N}\right\|)$ $\leq\frac{1}{N+1}\left(\bar{\kappa}+\bar{\kappa}\sum_{n=1}^{N}\frac{1}{n+1}+2\mu\sigma\sum_{n=1}^{N}\frac{1}{n}\right)\leq\frac{1}{N+1}\left(\bar{\kappa}+(\bar{\kappa}+2\mu\sigma)(\log{N}+1)\right),$
where $\bar{\kappa}=2\operatorname{dist}(x^{0},\mathop{\rm Fix}T)+\frac{\mu\sigma}{2}.$ Therefore the oracle complexity can be bounded by $\tilde{O}(\varepsilon^{-5})$. ∎
Remark 3.6.
There is a simple reason behind the choice of $k_{n}$ in this case. If one assumes that $k_{n}=n^{a}$, then $\sigma_{n}=O(n^{-a/2})$. We make $a\geq 2$ in order to verify condition ($H_{1}$). The use of (6) and simple computations leads to an oracle complexity bound of $\tilde{O}(\epsilon^{-h(a)})$, where
$h(a)=\begin{cases}\dfrac{2(a+1)}{a-2}&\text{ if }\,\,2<a\leq 4\\[5.69046pt] (1+a)&\text{ if }\,\,a\geq 4,\end{cases}$
which attains its minimum at $a=4$.
3.2. Lower bound
In this section, we obtain a lower bound for the number of queries required for any deterministic algorithm whose iterates $x^{n}$ are obtained as a linear combination of the previous iterates and the oracle approximation of $T$ at $x^{n-1}$ that may include variance reduction with minibatches. The construction of the operator is inspired by [38] and the notion of resistant oracle relies on ideas developed in [5].
Theorem 3.7.
Let $\sigma,\bar{\kappa}>0$, $(k_{n})_{n}\subseteq{\mathbb{N}}$ a sequence of minibatch sizes, and $0<\varepsilon<\frac{\sigma}{2}$ a given tolerance. Then, there exists a normed space $({\mathbb{R}}^{d},\left\|\cdot\right\|)$ (with $d$ depending on $\varepsilon$ and $\bar{\kappa}$), an operator $T$ (depending on $\varepsilon$), a stochastic oracle $\tilde{T}$, an initial point $x^{0}\in{\mathbb{R}}^{d}$ and a natural number $N=\Omega\left(\bar{\kappa}\sigma^{2}\varepsilon^{-3}\right)$ such that:
  • $T$ is nonexpansive and has bounded range $\bar{\kappa}$, i.e. $\left\|Tx\right\|\leq\bar{\kappa}$ for every $x\in{\mathbb{R}}^{d}$,
  • $\tilde{T}$ is unbiased and have uniformly bounded variance $\sigma^{2}$,
  • for any sequence $(x^{n})_{n}$ such that
    (7) $x^{n}\in\operatorname{span}\left\{x^{0},x^{1},...,x^{n-1},\frac{1}{k_{n}}\sum_{j=1}^{k_{n}}\tilde{T}(x^{n-1},\xi_{n,j})\right\},\quad\mbox{ for all }n\geq 1$
    we have
    ${\mathbb{E}}(\left\|x^{n}-Tx^{n}\right\|)>\varepsilon,\quad\mbox{for every $n$such that $\sum_{i=1}^{n}k_{i}\leq N$}$
In order to prove the Theorem, let us first define $\lambda=2\varepsilon$, and $d=\left\lfloor\frac{\bar{\kappa}}{\lambda}\right\rfloor$. We consider $\mathbb{R}^{d}$ endowed with the norm $\left\|x\right\|_{1}=\sum_{i=1}^{d}|x_{i}|$. Let the map $Q:{\mathbb{R}}^{d}\to{\mathbb{R}}^{d}$ be given by
$Q(x_{1},x_{2},\dots,x_{d})=(\lambda-x_{d},\;x_{1},x_{2},\dots,x_{d-1}).$
It is easy to check that $Q$ is nonexpansive (in fact, it is distance preserving). Moreover, $Q$ has a unique fixed-point $x^{*}=\left(\frac{\lambda}{2},\;\frac{\lambda}{2},\dots,\frac{\lambda}{2}\right)^{\top}\in\mathbb{R}^{d}$.
Let us define $C=\{x\in{\mathbb{R}}^{d}:\left\|x-x^{*}\right\|_{\infty}\leq\frac{\lambda}{2}\}$ the ball for the infinity norm centered on $x^{*}$ with radius $\lambda/2$. Let $P_{C}:{\mathbb{R}}^{d}\to C$ be the function defined by $(P_{C}(x))_{i}=P_{[0,\lambda]}(x_{i})$ for every $i=1,\dots,d$, where $P_{[0,\lambda]}$ is the one-dimensional projection onto the interval $[0,\lambda]$. Explicitly, $P_{[0,\lambda]}(t)$ equals $t$ if $t\in[0,\lambda]$, equals 00 if $t<0$, and equals $\lambda$ if $t>\lambda$.
Let $T:{\mathbb{R}}^{d}\to{\mathbb{R}}^{d}$ defined by $T=Q\circ P_{C}$. One can easily check that $T$ is nonexpansive for the norm $\left\|\cdot\right\|_{1}$, and $\mathop{\rm Fix}(T)=\{x^{*}\}$. Indeed,
$\left\|Tx-Ty\right\|_{1}=\left\|P_{C}(x)-P_{C}(y)\right\|_{1}=\sum_{i=1}^{d}|P_{[0,\lambda]}(x_{i})-P_{[0,\lambda]}(y_{i})|\leq\sum_{i=1}^{d}|x_{i}-y_{i}|.$
Here we have used the fact that $|P_{[0,\lambda]}(t)-P_{[0,\lambda]}(t^{\prime})|\leq|t-t^{\prime}|$ for all $t\in{\mathbb{R}}$.
Furthermore, for every $x\in{\mathbb{R}}^{d}$, we have $Tx\in C$. Indeed, let $y=P_{C}(x)$, then $y\in C$ and therefore
$\left\|Tx-x^{*}\right\|_{\infty}=\left\|Qy-x^{*}\right\|_{\infty}=\left\|y-x^{*}\right\|_{\infty}\leq\frac{\lambda}{2}.$
Using the triangular inequality we have
$\left\|Tx\right\|_{\infty}\leq\left\|Tx-x^{*}\right\|_{\infty}+\left\|x^{*}\right\|_{\infty}\leq\lambda,$
so $T$ has bounded range with $\left\|Tx\right\|_{1}\leq d\lambda\leq\bar{\kappa}$.
Now we will construct a stochastic oracle for $T$. For this, we define the progress of $x\in{\mathbb{R}}^{d}$ similarly to [5],
$\operatorname{prog}(x)=\max\{i\in\{1,\ldots,d\}\;:\;|x_{i}|>0\},\quad\text{and}\quad\operatorname{prog}(0_{{\mathbb{R}}^{d}})=0.$
Let $\xi$ be a random variable with a Bernoulli distribution of parameter $p\in(0,1)$. We define the stochastic oracle $\tilde{T}(x,\xi)$ component-wise as
$(\tilde{T}(x,\xi))_{i}=\begin{cases}(Tx)_{i}&\text{if}\quad i\neq\mbox{prog}(x)+1,\\[12.91663pt] \dfrac{\xi}{p}(Tx)_{i}&\text{if}\quad i=\mbox{prog}(x)+1,\end{cases}$
when $\mbox{prog}(x)<d$, and $\tilde{T}(x,\xi)=T(x)$ if $\mbox{prog}(x)=d$. It is easy to check that ${\mathbb{E}}(\tilde{T}(x,\xi))=Tx$ and
${\mathbb{E}}\left(\left\|Tx-\tilde{T}(x,\xi)\right\|^{2}_{2}\right)\leq((Tx)_{\operatorname{prog}(x)})^{2}{\mathbb{E}}\left[\left(1-\frac{\xi}{p}\right)^{2}\right]\leq\lambda^{2}\frac{1-p}{p},$
where in the last inequality we used that $\left\|Tx\right\|_{\infty}\leq\lambda$. Now, taking $p=\frac{\lambda^{2}}{\sigma^{2}}$ we have that ${\mathbb{E}}\left(||Tx-\tilde{T}(x,\xi)||^{2}_{2}\right)\leq\sigma^{2}$. Finally, the following lemma will be relevant to prove the lower bound. We leave its proof to Appendix A.
Lemma 3.8.
For all $x\in{\mathbb{R}}^{d}$ such that $\operatorname{prog}(x)<d$, we have $\left\|x-Tx\right\|_{1}\geq\lambda$.
Now we have collected all the ingredients to complete the proof.
Proof.
(of Theorem 3.7)
Let $N\in{\mathbb{N}}$ such that $N<\frac{d}{2p}\leq N+1$. Recalling the definitions of $\lambda$, $d$ and $p$, yields $N=\Omega(\kappa\sigma^{2}\varepsilon^{-3})$. We will prove that if $\sum_{i=1}^{n}k_{i}\leq N$ then ${\mathbb{E}}(\left\|x^{n}-Tx^{n}\right\|)>\varepsilon$.
Let $x^{0}=0_{\mathbb{R}^{d}}$, so $\text{prog}(x^{0})=0$. We observe that the stochastic oracle unveils a new coordinate in iteration $n$ if one or more of the samples $\xi_{n,1},\dots,\xi_{n,k_{n}}$ takes the value 1. Therefore, if $\text{prog}(x^{n-1})<d$, and $x^{n}$ were constructed according to (7), we have $\text{prog}(x^{n})=\text{prog}(x^{n-1})+1$ with probability $p_{n}=1-(1-p)^{k_{n}}$, and $\text{prog}(x^{n})=\text{prog}(x^{n-1})$ with probability $(1-p)^{k_{n}}$. As a result, the random variable $\text{prog}(x^{n})$ is the sum of $n$ independent Bernoulli random variables with parameters $p_{1},\dots,p_{n}$. In particular, the expected value of $\text{prog}(x^{n})$ is given by the sum of the probabilities. Now, using the Taylor expansion and Cauchy’s remainder, one can verify that $(1-p)^{k_{i}}\geq 1-pk_{i}$ for every $k_{i}\geq 1$. Hence, we have the estimate
$\sum_{i=1}^{n}p_{i}=\sum_{i=1}^{n}1-(1-p)^{k_{i}}\leq\sum_{i=1}^{n}pk_{i}\leq pN<\frac{d}{2}.$
Using Markov’s inequality, we have $\mathbb{P}(\mbox{prog}(x^{n})<d)\geq 1-\frac{\sum_{i=1}^{n}p_{i}}{d}>\frac{1}{2}$. Now, using Lemma 1 and Markov’s inequality for the random variable $\left\|x^{n}-Tx^{n}\right\|_{1}$ we have
$\frac{1}{2}<\mathbb{P}\left(\mbox{prog}(x^{n})<d\right)\leq\mathbb{P}\left(\left\|x^{n}-Tx^{n}\right\|_{1}\geq\lambda\right)\leq\frac{1}{\lambda}{\mathbb{E}}(\left\|x^{n}-Tx^{n}\right\|_{1}),$
and therefore $\varepsilon<{\mathbb{E}}(\left\|x^{n}-Tx^{n}\right\|_{1})$. ∎
Our lower bound depends on $\varepsilon$, $\bar{\kappa}$, and $\sigma$, all of which are also involved in the upper bound (6). Additionally, our results apply directly to both the Halpern iteration with minibatch discussed in this paper and the stochastic Krasnoselskii-Mann (sKM) iteration discussed in [15], which does not use variance reduction. In particular, this result reveals a gap of $\tilde{O}(\varepsilon^{-3})$ compared to the upper bound $\tilde{O}(\varepsilon^{-6})$ established for (sKM), as well as a gap of $\tilde{O}(\varepsilon^{-2})$ compared to the rate $\tilde{O}(\varepsilon^{-5})$ of the stochastic Halpern iteration. This prompts the consideration of improving the analysis or exploring alternative approaches to bridge the gap.
Remark 3.9.
In [16], the rate $\tilde{O}(\varepsilon^{-3})$ was achieved using the stochastic Halpern iteration within the Euclidean framework, specifically in the context of monotone inclusions. To do this, the authors require additional assumptions on $\tilde{T}$, as discussed in Section 2. Thus, our lower bound does not extend to such a specialized oracle. In fact, the shift operator constructed in this section does not satisfy the condition (3) and therefore the optimality of the approach in [16] remains an open question.
3.3. Convergence of the iterates
Even if we are mainly interested in establishing nonasymptotic bounds for the expected error term ${\mathbb{E}}(\left\|x^{n}-Tx^{n}\right\|)$, it is interesting to write an almost sure convergence result for the Halpern iteration (iH) for nonexpansive operators. The proof arguments are quite standard and rely on general results available for the noiseless iteration in Banach spaces (see [42]). We refer the reader to the technical details in the Appendix A.
Theorem 3.10.
Let $(x^{n})_{n}$ the sequence generated by (iH) where $T$ is nonexpansive, and $\sum_{n=1}^{\infty}\sigma_{n}<+\infty$. Assume also that the sequence $(\beta_{n})_{n}$ is such that
$\sum_{j=1}^{\infty}(1-\beta_{j})=+\infty,\quad\text{and}\quad\sum_{j=1}^{n}(\beta_{j}-\beta_{j-1})<+\infty.$
Then
  • $(a)$
    The fixed-point error $\left\|x^{n}-Tx^{n}\right\|$ goes to zero almost surely and any accumulation point of $(x^{n})_{n}$ is a fixed-point of $T$.
  • $(b)$
    If $\left({\mathbb{R}}^{d},\left\|\,\,\cdot\,\,\right\|\right)$ is a smooth space in the sense that the limit
    $\lim_{\lambda\to 0}\dfrac{\left\|x+\lambda\,y\right\|-\left\|x\right\|}{\lambda}$
    exists for all $x,y$ in the unit sphere, then $(x^{n})_{n}$ converges almost surely to some $x^{*}\in\mathop{\rm Fix}(T)$.
Remark 3.11.
Simple examples of smooth finite-dimensional Banach spaces covered by Theorem 3.10 are ${\mathbb{R}}^{d}$ endowed with the $\ell^{p}$ norm $\left\|x\right\|_{p}=(\sum_{i}|x_{i}|^{p})^{1/p}$ for $1<p<\infty$.
4. The contracting case
Let $T$ be a contraction with parameter $\gamma\in(0,1)$, where $x^{*}$ is the unique fixed-point of $T$. We are interested in the number of iterations and the overall oracle complexity required to obtain a solution $x^{n}$ such that ${\mathbb{E}}(\left\|x^{n}-x^{*}\right\|)\leq\varepsilon$ for a given tolerance $\varepsilon>0$.
Our recursive bound approach here is slightly different from the one in the nonexpansive case. We take increasing batch sizes that depend on the contraction parameter $\gamma$, allowing for a sensible decrease in the oracle complexity. The following result is the analogue of Proposition 3.1 in the contracting setting.
Proposition 4.1.
Let $T$ be a contraction with parameter $\gamma\in(0,1)$, $(\beta_{n})_{n}\subseteq(0,1)$ an increasing sequence of stepsiezes, and $(x^{n})_{n\geq 1}$ be the sequence generated by (iH). Then, for every $n\geq 1$, the following estimate holds
(8) $\left\|x^{n}-x^{*}\right\|\leq\sum_{i=0}^{n}\gamma^{n-i}B_{i+1}^{n}((1-\beta_{i})\left\|x^{0}-x^{*}\right\|+\beta_{i}\left\|U_{i}\right\|),$
Proof.
We proceed by induction. For $n=1$, one has that
$\left\|x^{1}\!-\!x^{*}\right\|$ $=\left\|(1-\beta_{1})(x^{0}-x^{*})+\beta_{1}(Tx^{0}-x^{*})+\beta_{1}U_{1}\right\|$
$\leq\beta_{1}\gamma\left\|x^{0}-x^{*}\right\|+(1-\beta_{1})\left\|x^{0}-x^{*}\right\|+\beta_{1}\left\|U_{1}\right\|$
Now, assuming that it holds for $n-1$, we have
$\left\|x^{n}\!-\!x^{*}\right\|$ $=\left\|(1-\beta_{n})(x^{0}-x^{*})+\beta_{n}(Tx^{n-1}-x^{*})+\beta_{n}U_{n}\right\|$
$\leq(1-\beta_{n})\left\|x^{0}-x^{*}\right\|+\gamma\beta_{n}\left\|x^{n-1}-x^{*}\right\|+\beta_{n}\left\|U_{n}\right\|$
$\leq(1-\beta_{1})\left\|x^{0}-x^{*}\right\|+\beta_{n}\left\|U_{n}\right\|+\gamma\beta_{n}\left(\sum_{i=0}^{n-1}\gamma^{n-1-i}B_{i+1}^{n-1}((1-\beta_{i})\left\|x^{0}-x^{*}\right\|+\beta_{i}\left\|U_{i}\right\|)\right)$
The result follows after reordering the last term. ∎
The next result makes explicit bounds for the expected fixed-point residual in the case where $\beta_{n}=\frac{n}{n+1}$ while shedding light about the appropriate minibatch size in this case.
Theorem 4.2.
Let $N\in{\mathbb{N}}$ be fixed and assume that $\beta_{n}=\frac{n}{n+1}$ and that
$\sigma_{n}^{2}\leq\frac{\sigma^{2}}{n^{2}\gamma^{N-n}},$
for some positive constant $\sigma^{2}$. Then,
${\mathbb{E}}(\left\|x^{N}-x^{*}\right\|)\leq\frac{\left\|x^{0}-x^{*}\right\|+2\sigma}{(1-\gamma)(N+1)}.$
Proof.
Taking expectation in (8) and using that $B_{i}^{n}=\frac{i}{n+1}$, we have
(9) ${\mathbb{E}}(\left\|x^{N}-x^{*}\right\|)$ $\leq\frac{1}{N+1}\sum_{i=0}^{N}\gamma^{N-i}(i+1)\left(\frac{1}{i+1}\left\|x^{0}-x^{*}\right\|+\frac{i}{i+1}\sigma_{i}\right)$
$\leq\frac{1}{N+1}\left(\frac{\left\|x^{0}-x^{*}\right\|}{1-\gamma}+\sigma\sum_{i=0}^{N}\gamma^{\frac{N-i}{2}}\right)$
$\leq\frac{1}{N+1}\left(\frac{\left\|x^{0}-x^{*}\right\|}{1-\gamma}+\frac{\sigma}{1-\sqrt{\gamma}}\right)$
The result follows by noticing that $\frac{1}{1-\sqrt{\gamma}}\leq\frac{2}{1-\gamma}$ for all $\gamma\in[0,1)$. ∎
The following corollary is the analogous of Corollary 3.5 as it covers the special case of uniformly bounded variance for contractive operators.
Corollary 4.3.
Let $\varepsilon>0$ be given, and assume that the stochastic oracle $\tilde{T}$ is with uniformly bounded variance with bound $\sigma^{2}$. Define $N=\left\lceil\frac{\left\|x^{0}-x^{*}\right\|+2\mu\sigma}{\varepsilon(1-\gamma)}\right\rceil$. Then Algorithm 1 with minibatches of size $k_{n}=\lceil n^{2}\gamma^{N-n}\rceil$ returns a solution $x^{N}$ such that ${\mathbb{E}}(\left\|x^{N}-x^{*}\right\|)\leq\varepsilon$, and it requires at most
$O\left(\frac{\left\|x^{0}-x^{*}\right\|^{2}+\mu^{2}\sigma^{2}}{\varepsilon^{2}(1-\gamma)^{3}}\right)$
queries to the stochastic oracle.
Proof.
The estimate (2) guarantees that $\sigma_{n}={\mathbb{E}}(\left\|U_{n}\right\|)\leq\mu\frac{\sigma}{\sqrt{k_{n}}}$. Therefore, for minibatches of size $k_{n}=\lceil n^{2}\gamma^{N-n}\rceil$, we have $\sigma_{n}\leq\mu\sigma(n\gamma^{\frac{N-n}{2}})^{-1}$.
Theorem 4.2 implies that for $N\geq\frac{\left\|x^{0}-x^{*}\right\|+2\mu\sigma}{\varepsilon(1-\gamma)}$ we have $\left\|x^{N}-x^{*}\right\|\leq\varepsilon$. Now, as every iteration $n=1,...,N$ requires $k_{n}$ queries to the stochastic oracle, we find an overall oracle complexity of
$\sum_{n=1}^{N}k_{n}\leq\sum_{n=1}^{N}n^{2}\gamma^{N-n}\leq 2\frac{N^{2}}{1-\gamma}+N\leq 16\frac{\left\|x^{0}-x^{*}\right\|^{2}+\mu^{2}\sigma^{2}}{(1-\gamma)^{3}\varepsilon^{2}}.$
Here, we have used $\lceil x\rceil\leq 2x$ for any $x\geq 1$. ∎
The overall oracle complexity described in the corollary relies on several factors: tolerance, contraction parameter, initial distance to the unique fixed-point, and variance of the noise, which might not be readily known beforehand. Nevertheless, in the case of $Q$-learning as explored later (c.f. Section 5.2), the critical parameters $\left\|x^{0}-x^{*}\right\|$ and $\sigma^{2}$ can be estimated straightforwardly. Consequently, one finds the more precise estimate $\tilde{O}(\varepsilon^{-2}(1-\gamma)^{-5})$.
5. Application to MDPs
Consider an infinite-horizon Markovian Decision Process (MDP) with state space $\mathcal{S}$ and action set $\mathcal{A}$, both assumed to be finite. Let $p(\cdot\,|\,s,a)\in\Delta(\mathcal{S})$ the probability distribution of the next state $s^{\prime}\in\mathcal{S}$ when one chooses the action $a\in\mathcal{A}$ in state $s\in\mathcal{S}$. The function $r:\mathcal{S}\times\mathcal{A}\to[0,1]$ is the reward function of the process, i.e. $r(s,a)$ is the reward obtained when the chain is in state $s$ and the action $a$ is chosen. Let $\pi:\mathcal{S}\to\Delta(\mathcal{A})$ be a possible randomized policy in the sense that $\pi(s)$ is the distribution of the action selected when the state is $s\in\mathcal{S}$. We say that a policy $\pi$ is deterministic if $\pi:\mathcal{S}\to\mathcal{A}$. Notice that the policies that we consider are stationary in the sense that they do not depend on time.
5.1. Average Reward Halpern $Q$-learning
In this part, we consider the so-called average reward setting, where one looks for a policy $\pi$ that maximizes the average reward
(10) $V_{\pi}(s)=\liminf_{N\to\infty}{\mathbb{E}}\left(\dfrac{1}{N}\sum_{k=0}^{N-1}r(s_{k},a_{k})\,\Big{|}\,s_{0}=s\right),\quad\text{for all}\,s\in\mathcal{S},$
where the trajectory $(s_{k},a_{k})_{k}$ is induced by $\pi$. It is well-known [40] that under the following unichain assumption
(uni) $&\text{for every stationary policy }\pi\text{ the induced Markov chain with transition}\\ &\text{probabilities }P_{ss^{\prime}}=p(s^{\prime}|s,\pi(s))\text{ has a single recurrent class,}\end{aligned}\right.$
[…]
$h(s)=\max_{a\in\mathcal{A}}\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s,a)\left(r(s,a)+h(s^{\prime})\right)-v^{*},\quad\forall\,s\in\mathcal{S}.$
To address the fact that transitions are unknown, one approach is to define the so-called $Q$-factors as
$Q(s,a)=\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s,a)\left(r(s,a)+h(s^{\prime})\right)-v^{*}.$
Hence, the previous equation can be written as
($Q_{\text{AV}}$) $Q(s,a)=\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s,a)(r(s,a)+\max_{a^{\prime}\in\mathcal{A}}Q(s^{\prime},a^{\prime}))-v^{*},\quad\forall\,(s,a)\in\mathcal{S}\times\mathcal{A}.$
Therefore the optimal $Q^{*}$ is the unique solution, up to a constant matrix, of the fixed-point equation $\mathcal{H}Q=Q$, where $\mathcal{H}:{\mathbb{R}}^{\mathcal{S}\times\mathcal{A}}\to{\mathbb{R}}^{\mathcal{S}\times\mathcal{A}}$ and, for all $(s,a)\in\mathcal{S}\times\mathcal{A}$,
(11) $\mathcal{H}Q(s,a):=\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s,a)\left(r(s,a)+\max_{a^{\prime}\in\mathcal{A}}Q(s^{\prime},a^{\prime})\right)-v^{*}.$
Consequently, the optimal policy $\pi^{*}$ can be obtained as $\pi^{*}(s)=\operatorname{argmax}_{a\in A}Q^{*}(s,a)$. It is well known that $\mathcal{H}$ is nonexpansive for the infinity norm $\left\|\,\cdot\,\right\|_{\infty}$. Although the problem of computing the $Q$-factors fits at first glance in our framework, one has to address the important issue related to the fact that the optimal value $v^{*}$ is unknown.
One of the few available methods in this case is the RVI-$Q$-learning algorithm, proposed in [1, Abounadi et al]. Explicitly, to overcome the problem of knowing $v^{*}$, the method introduces a function $f:{\mathbb{R}}^{\mathcal{S}\times\mathcal{A}}\to{\mathbb{R}}$ such that for every $c\in{\mathbb{R}}$, $f(Q+c\,{\bf e})=f(Q)+c$, where ${\bf e}\in{\mathbb{R}}^{\mathcal{S}\times\mathcal{A}}$ is the matrix full of ones. Typical examples of such functions are $f(Q)=\max_{s,a}Q(s,a)$, $f(Q)=\min_{s,a}Q(s,a)$, $f(Q)=\frac{1}{|\mathcal{S}\times\mathcal{A}|}\sum_{s,a}Q(s,a)$, and $f(Q)=Q(s_{0},a_{0})$ for some fixed $(s_{0},a_{0})\in\mathcal{S}\times\mathcal{A}$.
By taking an initial guess $Q^{0}$, for every $n\geq 1$ the update in RVI-$Q$-learning is
($Q_{\text{RVI}}$) $Q^{n}(s,a)=(1-\alpha_{n})Q^{n-1}(s,a)+\alpha_{n}\left(r(s,a)+\max_{a^{\prime}\in\mathcal{A}}Q^{n-1}(s_{n}(s,a),a^{\prime})-f(Q^{n-1})\right),$
where $s_{n}(s,a)$ is sampled from the distribution $p(\cdot|s,a)$ for every $(s,a)\in\mathcal{S}\times\mathcal{A}$ and $(\alpha_{n})_{n}$ is a sequence in $(0,1)$ such that $\sum_{n}\alpha_{n}=+\infty$ and $\sum_{n}\alpha_{n}^{2}<+\infty$. This is a synchronous algorithm in the sense that samples for every state-action pair $(s,a)$ are used at every iteration.
We stress that in the reinforcement learning literature, oracle complexity is measured considering not vector calls to the oracle but single calls to each component as a sample device. This means that our bounds will contain an extra $|\mathcal{S}\times\mathcal{A}|$ term that we make explicit.
Recently, in [15, Theorem 3.2] the authors established the first nonasymptotic rate for the expected value of the error residual in this case. Roughly speaking, if one takes a stepsize given by $\alpha_{n}=1/(n+1)^{a}$, $4/5<a\leq 1$, then ${\mathbb{E}}(\left\|Q^{n}-\mathcal{H}Q^{n}\right\|_{\infty})=O(|\mathcal{S}\times\mathcal{A}|^{1/2}n^{(1-a)/2})$. Then, in terms of oracle complexity, this means that, given a precision $\varepsilon>0$, the algorithm ($Q_{\text{RVI}}$) achieves a precision $\varepsilon$ after $O(|\mathcal{S}\times\mathcal{A}|^{3/2}\varepsilon^{-2/(1-a)})$ queries to the stochastic oracle since no variance reduction is used.
In what follows, we show how it is possible to improve the previous complexity by the use of the stochastic Halpern iteration in this case given by Algorithm 2 below is the main result of this section.
Algorithm 2 Halpern Synchronous $Q$-learning (Average reward)
1:Sequence $(\beta_{n})_{n}\in(0,1)$, initial point $Q^{0}$, function $f$ such that $f(Q+c\,{\bf e})=f(Q)+c$.
2:set $n\leftarrow 1$
3:for $n=1,\ldots,N$ do
4:     $k_{n}\leftarrow n^{6}$
5:     compute
~$\tilde{\mathcal{M}}_{n}Q^{n-1}(s,a)=r(s,a)+\frac{1}{k_{n}}\sum_{i=1}^{k_{n}}\max_{a^{\prime}\in\mathcal{A}}Q^{n-1}(s_{i}(s,a),a^{\prime})-f(Q^{n-1})$
where $s_{i}(s,a)\sim p(\cdot|s,a)$ for every $(s,a)\in\mathcal{S}\times\mathcal{A}$.
6:     set $Q^{n}=(1-\beta_{n})Q^{0}+\beta_{n}\tilde{\mathcal{M}}_{n}Q^{n-1}$
7:     $n\leftarrow n+1$
8:end for
9:Return $Q^{N}$
One challenge in analyzing Algorithm 2 is the fact that the underlying operator may not be nonexpansive, due to the choice of $f$. Consequently, we turn our attention to an alternative method, which we denote the Benchmark Algorithm 3 with state variable $Q_{v}^{n}$. It is important to note that Algorithm 3 is designed for theoretical analysis only and cannot be practically implemented.
Algorithm 3 Benchmark Halpern Synchronous $Q$-learning (Average reward)
1:Sequence $(\beta_{n})_{n}\in(0,1)$, initial point $Q^{0}$.
2:for $n=1,\ldots,N$ do
3:     $k_{n}\leftarrow n^{6}$
4:     compute
~$\tilde{\mathcal{H}}_{n}Q^{n-1}_{v}(s,a)=r(s,a)+\frac{1}{k_{n}}\sum_{i=1}^{k_{n}}\max_{a^{\prime}\in\mathcal{A}}Q^{n-1}_{v}(s_{i}(s,a),a^{\prime})-v^{*}$
where $s_{i}(s,a)\sim p(\cdot|s,a)$ for every $(s,a)\in\mathcal{S}\times\mathcal{A}$.
5:     set $Q^{n}_{v}=(1-\beta_{n})Q^{0}+\beta_{n}\tilde{\mathcal{H}}_{n}Q^{n-1}_{v}$
6:     $n\leftarrow n+1$
7:end for
8:Return $Q^{N}$
Let us consider the sequences $(Q^{n})_{n}$ and $(Q^{n}_{v})_{n}$ generated by Algorithms 2 and 3, respectively, using the same samples and initialized at a common point $Q^{0}$. We claim that for all $n\geq 0$, $Q^{n}-Q^{n}_{v}=c_{n}{\bf e}$ for some sequence of real numbers $(c_{n})_{n}$. Let us proceed by induction. The case $n=0$ is true by definition. Now, for every $(s,a)\in\mathcal{S}\times\mathcal{A}$ and using the induction hypothesis and the properties of the function $f$
$Q^{n}(s,a)\!-\!Q^{n}_{v}(s,a)$ $=\!\beta_{n}\left(\tilde{\mathcal{M}}_{n}Q^{n-1}(s,a)-\tilde{\mathcal{H}}_{n}Q^{n-1}_{v}(s,a))\right),$
$=\beta_{n}\left(\!\frac{1}{k_{n}}\sum_{i=1}^{k_{n}}\max_{a^{\prime}\in\mathcal{A}}Q^{n-1}(s_{i}(s,a),a^{\prime})\!-\!f(Q^{n-1})\!-\!\frac{1}{k_{n}}\sum_{i=1}^{k_{n}}\max_{a^{\prime}\in\mathcal{A}}Q^{n-1}_{v}(s_{i}(s,a),a^{\prime})+v^{*}\!\right),$
$=\beta_{n}\left(\frac{1}{k_{n}}\sum_{i=1}^{k_{n}}\max_{a^{\prime}\in\mathcal{A}}Q_{v}^{n-1}(s_{i}(s,a),a^{\prime})\!+\!c_{n-1}\!-\!f(Q_{v}^{n-1}\!\!-\!\!c_{n-1}{\mathbf{e}})\right.$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\left.-\frac{1}{k_{n}}\sum_{i=1}^{k_{n}}\max_{a^{\prime}\in\mathcal{A}}Q^{n-1}_{v}(s_{i}(s,a),a^{\prime})\!+\!v^{*}\right),$
$=\beta_{n}\left(v^{*}-f(Q^{n-1}_{v})\right),$
which is independent of $(s,a)$. In particular, we have ${\mathbb{E}}(\left\|\mathcal{H}Q^{n}-Q^{n}\right\|_{\infty})={\mathbb{E}}(\left\|\mathcal{H}Q^{n}_{v}-Q^{n}_{v}\right\|_{\infty})$.
The following result establishes the convergence rate for ${\mathbb{E}}(\left\|\mathcal{H}Q^{n}-Q^{n}\right\|_{\infty})$ in the case where $\beta_{n}=\frac{n}{n+1}$. As a consequence of the previous argument, it is sufficient to obtain a convergence rate for the benchmark sequence ${\mathbb{E}}(\left\|\mathcal{H}Q^{n}_{v}-Q^{n}_{v}\right\|_{\infty})$.
Theorem 5.1.
For $\beta_{n}=\frac{n}{n+1}$ and a given precision $\varepsilon>0$ the sequence $(Q^{n})_{n}$ generated by Algorithm 2 produces an iterate $Q^{N}$ such that ${\mathbb{E}}\left(\left\|\mathcal{H}Q^{N}-Q^{N}\right\|\right)\leq\varepsilon$ with a number of calls to the sampling device of at most $\tilde{O}\left(|\mathcal{S}\times\mathcal{A}|\,\varepsilon^{-7}\right)$.
Proof.
Let us analyze the benchmark Algorithm 3. Set $g=r_{\max}-v^{*}$, where $r_{\max}:=\max_{(s,a)}r(s,a)$. Let us observe that
$\left\|U_{n}\right\|_{\infty}=\left\|\tilde{\mathcal{H}}_{n}Q^{n-1}_{v}-\mathcal{H}Q^{n-1}_{v}\right\|_{\infty}\leq 2\left\|Q_{v}^{n-1}\right\|_{\infty},$
where $\tilde{\mathcal{H}}$ is explicitly defined in Algorithm 3. Note that each component of $U_{n}$ is the sum of $k_{n}$ i.i.d. random variables with zero mean and bounded in absolute value by $\frac{2\left\|Q^{n-1}_{v}\right\|}{k_{n}}$. Let $\delta\in(0,1)$. Using Hoeffding’s inequality we have that
$\left\|U_{n}\right\|_{\infty}\leq\frac{\left\|Q^{n-1}_{v}\right\|_{\infty}\sqrt{8\log\left(\frac{2|\mathcal{S}\times\mathcal{A}|}{\delta}\right)}}{\sqrt{k_{n}}},\qquad\mbox{w.p. at least $1-\delta$}.$
Using the two bounds above we have
(12) ${\mathbb{E}}(\left\|U_{n}\right\|_{\infty})\leq(1-\delta)\frac{\left\|Q^{n-1}_{v}\right\|_{\infty}\sqrt{8\log\left(\frac{2|\mathcal{S}\times\mathcal{A}|}{\delta}\right)}}{\sqrt{k_{n}}}+2\delta\left\|Q^{n-1}_{v}\right\|_{\infty}$
On the other hand, from the definition of the iteration it follows that
$\left\|Q_{v}^{n}\right\|_{\infty}=\left\|(1-\beta_{n})Q^{0}+\beta_{n}\tilde{\mathcal{H}}_{n}Q_{v}^{n-1}\right\|_{\infty}\leq(1-\beta_{n})\left\|Q^{0}\right\|_{\infty}+\beta_{n}(g+\left\|Q_{v}^{n-1}\right\|_{\infty})$
Therefore, summing over the previous bound and using the precise definition of $\beta_{n}$, we obtain the estimation
$\left\|Q_{v}^{n}\right\|_{\infty}$ $\leq\frac{1}{n+1}\left\|Q^{0}\right\|_{\infty}\!\!+\!\frac{1}{n+1}\sum_{i=1}^{n}(i+1)\left(\frac{1}{i+1}\left\|Q^{0}\right\|_{\infty}+\frac{i}{i+1}g\right)$
(13) $=\left\|Q^{0}\right\|_{\infty}+\frac{n}{2}g$
This shows that one can obtain an a priori limit on the rate with which the state variable can increase. Now, recalling that $\sigma_{n}={\mathbb{E}}(\left\|U_{n}\right\|)$ and using (13), $k_{n}=n^{6}$, and $\delta=\frac{1}{\sqrt{k_{n}}}$ in (12) we obtain the estimate
$n\sigma_{n}\leq\frac{\left\|Q^{0}\right\|_{\infty}+\frac{n-1}{2}g}{n^{2}}\left(\sqrt{8\log\left(2n^{3}|\mathcal{S}\times\mathcal{A}|\right)}+2\right),$
and therefore $\sum_{n=1}^{N}n\sigma_{n}=\tilde{O}(1)$. Using Theorem 3.3, we find a solution $Q^{N}_{v}$ such that ${\mathbb{E}}(\left\|{\mathcal{H}}Q^{N}_{v}-Q^{N}_{v}\right\|_{\infty})\leq\varepsilon$ after $N=\tilde{O}(1/\varepsilon)$ iterations. As each iteration $n=1,...,N$ uses a total of $|\mathcal{S}\times\mathcal{A}|n^{6}$ queries, we obtain an overall oracle complexity of
$O(|\mathcal{S}\times\mathcal{A}|N^{7})=\tilde{O}(|\mathcal{S}\times\mathcal{A}|\varepsilon^{-7}).$
5.2. Discounted Halpern $Q$-learning
In this part, we consider the so-called discounted setting, where we look for a policy $\pi$ that maximizes the discounted reward
$V_{\pi}(s)={\mathbb{E}}\left(\sum_{k=0}^{\infty}\gamma^{k}r(s_{k},a_{k})\,\Big{|}\,s_{0}=s\right),\quad\text{for all}\,s\in\mathcal{S}.$
The optimal value function $V_{*}:\mathcal{S}\to{\mathbb{R}}$ is defined as $V_{*}(s)=\max_{\pi}V_{\pi}(s)$. It is well known that there exists an optimal deterministic policy $\pi_{*}$ such that $V_{*}=V_{\pi_{*}}$ and the optimal value function $V_{*}$ is the unique solution to the Bellman equations
$V(s)=\max_{a\in\,A}\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s,a)\left(r(s,a)+\gamma V(s^{\prime})\right),\quad\forall\,s\in\mathcal{S}.$
Therefore, the $Q$-factors in this case are
$Q(s,a)=\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s,a)\left(r(s,a)+\gamma V(s^{\prime})\right),$
and then
$Q(s,a)=\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s,a)\left(r(s,a)+\gamma\max_{a^{\prime}\in\mathcal{A}}Q(s^{\prime},a^{\prime})\right),\quad\forall\,(s,a)\in\mathcal{S}\times\mathcal{A}.$
In particular, the optimal $Q^{*}$ is the unique solution of the fixed-point equation $\mathcal{T}Q=Q$, where $\mathcal{T}:{\mathbb{R}}^{\mathcal{S}\times\mathcal{A}}\to{\mathbb{R}}^{\mathcal{S}\times\mathcal{A}}$ is the right-hand side of the last equation which is $\gamma$-contracting for the infinity norm.
One of the most studied methods is the so-called vanilla version $Q$-learning algorithm [53]. Explicitly, the method chooses $\left\|Q^{0}\right\|_{\infty}\leq\frac{r_{max}}{1-\gamma}$ and, for every $n\geq 1$ the update is
($Q_{\gamma}$) $Q^{n}(s,a)=(1-\alpha_{n})Q^{n-1}(s,a)+\alpha_{n}\left(r(s,a)+\gamma\max_{a^{\prime}\in\mathcal{A}}Q^{n-1}(s_{n}(s,a),a^{\prime})\right),$
where $s_{n}(s,a)$ is sampled from the distribution $p(\cdot|s,a)$ for every $(s,a)\in\mathcal{S}\times\mathcal{A}$ and $(\alpha_{n})_{n}\subseteq(0,1)$ is the sequence of learning rates. As we already mentioned, this is one of the most studied methods in the reinforcement learning literature.
It is worth noting that in this case the oracle has uniformly bounded variance; therefore, it is possible to directly apply Corollary 4.3 in order to obtain an oracle complexity bound. Indeed, using the bound (14) below, one can estimate the term $\sigma^{2}$ precisely for each coordinate. However, when bounding the variance in the Euclidean, the estimate carries an additional $\sqrt{|\mathcal{S}\times\mathcal{A}|}$ term. Then, using also that $\mu=1$ and that $\left\|Q^{0}-Q^{*}\right\|\leq\frac{\gamma r_{\max}}{1-\gamma}$ the complexity achieved by the mentioned corollary is
$O\left(\frac{|\mathcal{S}\times\mathcal{A}|^{2}r_{\max}^{2}}{\varepsilon^{2}(1-\gamma)^{5}}\right),$
which is worst than the rate claimed in Theorem  5.2 below.
Algorithm 4 corresponds to the stochastic Halpern Algorithm 1 in this precise context.
Algorithm 4 Halpern Synchronous $Q$-learning (discounted case)
1:Sequence $(\beta_{n})_{n}\in(0,1)$, initial point $Q^{0}$, discount factor $\gamma$, number of iterations $N$.
2:set $n\leftarrow 1$
3:for $n=1,\ldots,N$ do
4:     $k_{n}\leftarrow\lceil n^{2}\gamma^{N-n}\rceil$
5:     compute
$\hat{\mathcal{T}}_{n}Q^{n-1}(s,a)=r(s,a)+\gamma\frac{1}{k_{n}}\sum_{i=1}^{k_{n}}\max_{a^{\prime}\in\mathcal{A}}Q^{n-1}(s_{i}(s,a),a^{\prime})$
where $s_{i}(s,a)\sim p(\cdot|s,a)$ for every $(s,a)\in\mathcal{S}\times\mathcal{A}$.
6:     set $Q^{n}=(1-\beta_{n})Q^{0}+\beta_{n}\hat{\mathcal{T}}_{n}Q^{n-1}$
7:     $n\leftarrow n+1$
8:end for
9:Return $Q^{N}$
Theorem 5.2.
Assume that the initial condition $Q^{0}$ is such that $\left\|Q^{0}\right\|_{\infty}\leq\frac{r_{\max}}{1-\gamma}$. Then, given $\epsilon>0$ the iteration generated by Algorithm 4 with $\beta_{n}=\frac{n}{n+1}$ produces an iterate $Q^{N}$ such that
${\mathbb{E}}\left(\left\|Q^{N}-Q^{*}\right\|_{\infty}\right)\leq\varepsilon,$
using at most $\tilde{O}\left(\frac{|\mathcal{S}\times\mathcal{A}|r_{\max}^{2}}{\varepsilon^{2}(1-\gamma)^{5}}\right)$ queries to the stochastic oracle.
Proof.
A simple induction shows that if $\left\|Q^{0}\right\|_{\infty}\leq\frac{r_{\max}}{1-\gamma}$, then $\left\|Q^{n}\right\|_{\infty}\leq\frac{r_{\max}}{1-\gamma}$ for all $n\geq 1$. Now, we have the estimate
(14) $\left\|U_{n}\right\|_{\infty}\leq\left\|\hat{\mathcal{T}}_{n}\,Q^{n-1}-\mathcal{T}\,Q^{n-1}\right\|_{\infty}\leq 2\gamma\left\|Q^{n-1}\right\|_{\infty}\leq\frac{2\gamma r_{\max}}{1-\gamma}.$
Similarly to the average-reward case, we observe that the components of $U_{n}$ are the sum of $k_{n}$ i.i.d. random variables with zero mean and bounded in absolute value by $\frac{2\gamma r_{\max}}{k_{n}(1-\gamma)}$. Using Hoeffding’s inequality we obtain that
$\left\|U_{n}\right\|_{\infty}\leq\frac{\gamma r_{\max}\sqrt{8\log\left(\frac{2|\mathcal{S}\times\mathcal{A}|}{\delta}\right)}}{\sqrt{k_{n}}(1-\gamma)},\qquad\mbox{w.p. at least $1-\delta$}.$
Using the estimate (14) we have
(15) ${\mathbb{E}}(\left\|U_{n}\right\|_{\infty})\leq(1-\delta)\frac{\gamma r_{\max}\sqrt{8\log\left(\frac{2|\mathcal{S}\times\mathcal{A}|}{\delta}\right)}}{\sqrt{k_{n}}(1-\gamma)}+\delta\frac{2\gamma r_{\max}}{(1-\gamma)}$
Hence, taking $\delta=\frac{1}{\sqrt{k_{n}}}$ and $k_{n}=\lceil n^{2}\gamma^{N-n}\rceil$, we have the estimate
$\sigma_{n}={\mathbb{E}}(\left\|U_{n}\right\|_{\infty})=\tilde{O}\left(\frac{\gamma r_{\max}}{(1-\gamma)n\gamma^{\frac{N-n}{2}}}\right).$
Recalling equation (9) and proceeding similarly to the proof of Proposition 4.1 we find the rate
${\mathbb{E}}(\left\|Q^{N}-Q^{*}\right\|_{\infty})=\tilde{O}\left(\frac{\left\|Q^{0}-Q^{*}\right\|_{\infty}+\frac{r_{\max}}{(1-\gamma)}}{N(1-\gamma)}\right)=\tilde{O}\left(\frac{r_{\max}}{N(1-\gamma)^{2}}\right),$
where we have used that $\left\|Q^{*}\right\|_{\infty}\leq\frac{r_{\max}}{(1-\gamma)}$ (see [49, Lemma 1]).
Therefore, to reach the precision $\varepsilon$ we need at most $N=\tilde{O}\left(\frac{r_{\max}}{\varepsilon(1-\gamma)^{2}}\right)$ iterations, and the total number of samples is
$|\mathcal{S}\times\mathcal{A}|\sum_{n=1}^{N}k_{n}=|\mathcal{S}\times\mathcal{A}|\sum_{n=1}^{N}n^{2}\gamma^{N-n}=O\left(\frac{|\mathcal{S}\times\mathcal{A}|N^{2}}{(1-\gamma)}\right)=\tilde{O}\left(\frac{|\mathcal{S}\times\mathcal{A}|r_{\max}^{2}}{\varepsilon^{2}(1-\gamma)^{5}}\right).$
Interestingly, Theorem 5.2 recovers the best rate known until recently. This rate was subsequently improved to match, up to log terms, the complexity $O((1-\gamma)^{-3}\varepsilon^{-2})$ of the model-based version [6] by [50] through variance reduction.
Also, in the same paper, the authors showed that incorporating a restart strategy along with minibatches results in an improved rate in terms of $\frac{1}{1-\gamma}$. It would be interesting to investigate whether these concepts can be applied to the Halpern iteration. We defer this exploration to future research.
Appendix A Deferred proofs
Proof of Theorem 3.10
$(a)$ From (5), we have that
$\left\|x^{n}-x^{*}\right\|\leq\left\|x^{0}-x^{*}\right\|+\sum_{i=1}^{n}B^{n}_{i}\left\|U_{i}\right\|\leq\left\|x^{0}-x^{*}\right\|+\sum_{i=1}^{n}\left\|U_{i}\right\|.$
Using the monotone convergence theorem, the random variable $\sum_{i=1}^{n}\left\|U_{i}\right\|$ converges almost surely to the positive and integrable random variable $\sum_{i=1}^{\infty}\left\|U_{i}\right\|$, since $\sum_{n=1}^{\infty}\sigma_{n}<\infty$. Therefore, $(x^{n})_{n}$ is almost surely bounded and we conclude that $\kappa:=\sup_{n\geq 1}\left\|Tx^{n}-x^{0}\right\|$ is a.s. bounded.
Therefore, from (4), almost surely
$\left\|x^{n}-Tx^{n}\right\|\leq\kappa(1-\beta_{n})+\sum_{i=1}^{n}B_{i}^{n}\left(\kappa(\beta_{i}-\beta_{i-1})+\beta_{i}\left\|U_{i}\right\|+\beta_{i-1}\left\|U_{i-1}\right\|\right)+\beta_{n}\left\|U_{n}\right\|.$
We aim to prove that each of the terms involved goes to zero almost surely. First, given that $\lim_{n\to\infty}\beta_{n}=1$ one has that $\kappa(1-\beta_{n})$ vanishes. Also, $\lim_{n\to\infty}\left\|U_{n}\right\|=0$ a.s. Let us now fix a trajectory where $\sum_{i=1}^{\infty}\left\|U_{i}\right\|<+\infty$.
Observe that for every $i\in{\mathbb{N}}$, using the classical exponential bound, we have that
$B_{i}^{n}=\prod_{j=i}^{n}\beta_{j}\leq\exp(-\sum_{j=i}^{n}(1-\beta_{j})).$
[…]
To finish the proof, it suffices to note that since $(x^{n})_{n}$ is bounded and $\left\|x^{n}-Tx^{n}\right\|$ goes to zero, any cluster point belongs to $\mathop{\rm Fix}T$.
$(b)$ This part follows closely the arguments used by [54, Theorem 3.1] to study strong convergence in the noiseless case for general Banach spaces. For $10<t<$ 1, let $z^{t}$ the unique solution of the fixed-point equation
(16) $z^{t}=(1-t)\,x^{0}+t\,Tz^{t}.$
A direct application [42, Corollary 1] shows that $z^{*}:=\lim_{t\to 1}z^{t}$ exists and is a fixed-point of $T$. Let $J:({\mathbb{R}}^{d},\left\|\cdot\right\|)\to({\mathbb{R}}^{d},\left\|\cdot\right\|_{*})$ be the (standard) duality map defined as the subdifferential of the function $\frac{1}{2}\left\|\cdot\right\|^{2}$, i.e.
$J(x)=\{x^{*}\in{\mathbb{R}}^{d}\,|\,\langle x,x^{*}\rangle=\left\|x\right\|^{2}=\left\|x^{*}\right\|_{*}^{2}\},$
[…]
It is well known that a smooth finite-dimensional Banach space is uniformly smooth, and then $J$ is single valued and continuous on bounded sets. Using the subdifferential inequality and following exactly the same lines as in the proof of [54, Theorem 3.1], we obtain that, for all $n\geq 1$,
$\left\|x^{n}-z^{t}\right\|\leq$ $(1+(1-t)^{2})\left\|z^{t}-x^{n}\right\|^{2}+\left\|x^{n}-Tx^{n}\right\|\left(2\left\|z^{t}-x^{n}\right\|+\left\|x^{n}-Tx^{n}\right\|\right)+$
$+2(1-t)\langle x^{0}-z^{t},J(x^{n}-z^{t})\rangle,$
and then,
$\langle x^{0}-z^{t},J(z^{t}-x^{n})\rangle\leq\frac{(1-t)}{2}\left\|z^{t}-x^{n}\right\|^{2}+\frac{\left\|x^{n}-Tx^{n}\right\|}{2(1-t)}\left(2\left\|z^{t}-x^{n}\right\|+\left\|x^{n}-Tx^{n}\right\|\right).$
Using that $\left\|x^{n}-Tx^{n}\right\|$ goes to zero almost surely one has that a.s.
$\limsup_{n\to+\infty}\langle x^{0}-z^{t},J(z^{t}-x^{n})\rangle\leq\limsup_{n\to+\infty}\frac{(1-t)}{2}\left\|z^{t}-x^{n}\right\|^{2},$
and, since $x^{n}$ is bounded and $J$ is continuous, taking the limit when $t\to 1$ one has that
(17) $\limsup_{n\to+\infty}\langle x^{0}-z^{*},J(z^{*}-x^{n})\rangle\leq 0.$
Now, on the other hand, using again the subdifferential inequality
$\left\|x^{n}-z^{*}\right\|^{2}=$ $\left\|(1-\beta_{n})(x^{0}-z^{*})+\beta_{n}(Tx^{n-1}-z^{*}+U_{n})\right\|^{2}$
$\leq$ $\beta_{n}^{2}\left\|Tx^{n-1}-z^{*}+U_{n}\right\|^{2}+2(1-\beta_{n})^{2}\langle x^{0}-z^{*},J(x^{n}-z^{*})\rangle,$
$\leq$ $\beta_{n}\left\|x^{n-1}-z^{*}\right\|^{2}+\beta_{n}^{2}\left(2\left\|x^{n-1}-z^{*}\right\|\left\|U_{n}\right\|+\left\|U_{n}\right\|^{2}\right)$
$\hskip 64.58313pt+2(1-\beta_{n})\langle x^{0}-z^{*},J(x^{n}-z^{*})\rangle.$
Finally, noticing that $\sum_{n\geq 1}\beta_{n}^{2}\left(2\left\|x^{n-1}-z^{*}\right\|\left\|U_{n}\right\|+\left\|U_{n}\right\|^{2}\right)$ is a.s. finite along with (17) one can readily use [54, Lemma 2.5] to conclude that $\left\|x_{n}-z^{*}\right\|$ goes to zero almost surely.
Proof of Lemma 3.8
For $n<d$, let us define $\phi_{n}:{\mathbb{R}}^{d}\to{\mathbb{R}}$ for
$\phi_{n}(x)=|\lambda-x_{1}|+\sum_{i=2}^{n}|P_{[0,\lambda]}(x_{i-1})-x_{i}|+P_{[0,\lambda]}(x_{n}).$
We will prove by induction over $n$ that $\phi_{n}(x)\geq\lambda$ for every $x\in{\mathbb{R}}^{d}$. The result of the lemma follows by noticing that if $\mbox{prog}(x)<d$ then $\left\|x-Tx\right\|_{1}=\phi_{\mbox{prog}(x)}(x)$.
For $n=1$, $\phi_{1}(x)=|\lambda-x_{1}|+P_{[0,\lambda]}(x_{1})$. We have three cases:
  • if $x_{1}\in[0,1]$ then $\phi_{1}(x_{1})=(\lambda-x_{1})+x_{1}=\lambda$,
  • if $x_{1}>\lambda$ then $\phi_{1}(x_{1})=(x_{1}-\lambda)+\lambda=x_{1}>\lambda$,
  • if $x_{1}<0$ then $\phi_{1}(x_{1})=(\lambda-x_{1})>\lambda$,
hence, the results is valid for $n=1$.
Assume that $\phi_{n-1}(x)\geq\lambda$ for all $x\in{\mathbb{R}}^{d}$. We can write
$\phi_{n}(x)$ $=|\lambda-x_{1}|+\sum_{i=1}^{n}|P_{[0,\lambda]}(x_{i-1})-x_{i}|+P_{[0,\lambda]}(x_{n})$
$=\phi_{n-1}(x)+|P_{[0,\lambda]}(x_{n-1})-x_{n}|+P_{[0,\lambda]}(x_{n})-P_{[0,\lambda]}(x_{n-1})$
$\geq\lambda+|P_{[0,\lambda]}(x_{n-1})-x_{n}|+P_{[0,\lambda]}(x_{n})-P_{[0,\lambda]}(x_{n-1})$
Therefore, it is sufficient to prove that the function $\eta(z,z^{\prime}):=|P_{[0,\lambda]}(z)-z^{\prime}|+P_{[0,\lambda]}(z^{\prime})-P_{[0,\lambda]}(z)$ is nonnegative for all $(z,z^{\prime})\in{\mathbb{R}}^{2}$. We analyze the following cases:
  • if $z\in[0,\lambda]$ then for the nonexpansiveness of $P_{[0,\lambda]}$ we have
    $P_{[0,\lambda]}(z)-P_{[0,\lambda]}(z^{\prime})\leq|P_{[0,\lambda]}(z)-P_{[0,\lambda]}(z^{\prime})|\leq|z-z^{\prime}|=|P_{[0,\lambda]}(z)-z^{\prime}|,$
    which implies $\eta(z,z^{\prime})\geq 0$ for all $z^{\prime}\in{\mathbb{R}}$.
  • If $z>\lambda$ then $\eta(z,z^{\prime})=|\lambda-z^{\prime}|+P_{[0,\lambda]}(z^{\prime})-\lambda$ which is nonnegative again for the nonexapnsiveness of $P_{[0,\lambda]}$, namely
    $\lambda-P_{[0,\lambda]}(z^{\prime})=|P_{[0,\lambda]}(\lambda)-P_{[0,\lambda]}(z^{\prime})|\leq|\lambda-z^{\prime}|.$
  • Finally, if $z<0$ we have $\eta(z,z^{\prime})=|z^{\prime}|+P_{[0,\lambda]}(z^{\prime})$ which is clearly nonnegative.